# Scaling Relations in the Phase-Space Structure of Dark Matter Haloes

### v1

@@ Abstract
pubyear: 2024pagerange: Scaling Relations in the Phase-Space Structure of Dark Matter HaloesReferences
@@ 1 Introduction
Other scaling relations for simulated haloes have been reported in the literature. Notably, as first presented by Taylor & Navarro (2001) and subsequently explored by many others (e.g., Austin et al. 2005; Hoffman et al. 2007; Ludlow et al. 2011; Arora & Williams 2020), the pseudo-phase-space density follows a power law, $\rho(r)/\sigma^{3}(r)\propto r^{-\alpha}$ with $\alpha=1.875$, where $\rho(r)$ and $\sigma(r)$ are the density and velocity dispersion at radius $r$, respectively. In addition, the volumetric density of the DF, ${\cal{V}}(f_{0})=\int d^{3}\bm{r}d^{3}\bm{v}\delta(f(\bm{r},\bm{v})-f_{0})$, was shown to approximately follow the power law ${\cal{V}}(f)\propto f^{-2.5}$ over a wide range of $f$ values (Arad et al., 2004). We will show that this result on ${\cal{V}}(f)$ can be accounted for by our scaling relations.
@@ 3 Scaling Relations for Simulated Haloes
Motivated by the SIS scaling relations, we now present similar scaling relations for simulated haloes. It is well known that the density profile of dark matter haloes can be well fitted by the NFW profile (Navarro et al., 1997) between $0.05R_{\mathrm{vir}}$ and $R_{\mathrm{vir}}$ with $R_{\mathrm{vir}}$ being the virial radius, though deviations can occur outside this range (Schaller et al., 2015). The NFW profile is given by
@@
where $c=R_{\mathrm{vir}}/r_{\mathrm{s}}$ is the halo concentration, $\alpha$ and $\beta$ are dimensionless constants, and $m$ and $n$ are power indices. The NFW profile changes from $\rho\propto r^{-1}$ at $r\ll r_{\mathrm{s}}$ to $\rho\propto r^{-3}$ at $r\gg r_{\mathrm{s}}$, and behaves like the SIS at $r\sim r_{\mathrm{s}}$. While the SIS scaling relations suggest $m\approx 1$ and $n\approx-2$, we will obtain the best-fit values for $m$ and $n$ along with those for $\alpha$ and $\beta$ from the data on simulated haloes.
@@ 3.1 Halo Sample
As in Gross et al. (2024), we use the haloes in the TNG300-1-Dark simulations from the IllustrisTNG Project (Springel et al., 2017; Pillepich et al., 2017; Nelson et al., 2017; Naiman et al., 2018; Marinacci et al., 2018). The simulations were performed with the Planck Collaboration XIII (Planck Collaboration et al., 2016) cosmological parameters: $\Omega_{\mathrm{m}}=0.3089$, $\Omega_{\mathrm{b}}=0.0486$, $\Omega_{\Lambda}=0.6911$, $h=0.6774$, $n_{\mathrm{s}}=0.9667$, and $\sigma_{8}=0.8159$. The particle mass is $m_{\rm d}=7\times 10^{7}h^{-1}M_{\odot}$ and the softening length is $1h^{-1}\mathrm{kpc}$. The virial mass and radius of a halo, $M_{\mathrm{vir}}$ and $R_{\mathrm{vir}}$, respectively, are determined such that the average density inside $R_{\mathrm{vir}}$ is equal to $\Delta_{\mathrm{vir}}$ times the critical density of the Universe, where $\Delta_{\mathrm{vir}}=18\pi^{2}+82(\Omega_{\mathrm{m}}-1)-39(\Omega_{\mathrm{m}}-1)^{2}\approx 102$ is the virial factor (Bryan & Norman, 1998).
@@
For each halo of the final sample, we construct $dM/dE$ by counting the particles within $r_{\textrm{lim}}=2.5R_{\mathrm{vir}}$, which is approximately the depletion radius separating a growing halo from its draining environment (Fong & Han, 2021). As the energy is very sensitive to the choice of reference frame, we use the mean values for the most bound 10% of the particles in a halo to define its position and velocity (cf. Han et al., 2012). The DF is obtained as $f(E)=(dM/dE)/g(E)$ (e.g., Natarajan et al. 1997). Because we only count particles with $r<r_{\textrm{lim}}$, the corresponding $g(E)$ is
@@ 3.2 Fitting of Scaling Relations
Rather than fitting the scaling relation for $dM/dE$ in Eq. (17) to each halo, we aim to find a relation that applies to all the haloes. We obtain the numerical constant $\alpha$ and power index $m$ in this relation as follows. Because particles with $r\sim R_{\mathrm{vir}}$ are of main concern, we use a uniform logarithmic grid of $r_{\mathrm{max}}(E)/R_{\mathrm{vir}}$ for all the haloes. For each halo, we recalculate $dM/dE$ by counting the particles in each logarithmic bin of $r_{\mathrm{max}}(E)/R_{\mathrm{vir}}$ (i.e., $E$ is an implicit variable relating $r_{\mathrm{max}}(E)$ to $dM/dE$). Note that over the fitting range below, there are at least 10 particles in each bin for every halo so that $dM/dE$ can be calculated to reasonable accuracy. For each bin, we sort values of $(dM/dE)Gc^{1-m}/R_{\mathrm{vir}}$ for all the haloes to find the median. By inspection, the median value of $(dM/dE)Gc^{1-m}/R_{\mathrm{vir}}$ as a function of $r_{\mathrm{max}}(E)/R_{\mathrm{vir}}$ follows the scaling relation in Eq. (17) over $r_{\mathrm{max}}(E)\in[0.02R_{\mathrm{vir}},R_{\mathrm{vir}}]$. Detailed fitting by minimizing the rms deviation of $\log[(dM/dE)Gc^{1-m}/R_{\mathrm{vir}}]$ over this range gives $\alpha=0.60$ and $m=1.01$ with $\delta_{\log[(dM/dE)Gr_{\mathrm{s}}^{0.01}/R_{\mathrm{vir}}^{1.01}]}=0.019$. Because the fitted $m$ value is so close to 1, we repeat the above procedure by setting $m=1$ and obtain $\alpha=0.60$ with $\delta_{\log[(dM/dE)G/R_{\mathrm{vir}}]}=0.020$. In view of the simplicity without loss of accuracy, we adopt
@@
With $\delta_{\log[(dM/dE)G/R_{\mathrm{vir}}]}=0.019$ and $\delta_{\log[f(E)Gv_{\mathrm{s}}R_{\mathrm{vir}}^{2}]}=0.011$, these alternate scaling relations also provide good description of the halo data, but will not be discussed further in view of their similarity to the scaling relations adopted above. For our discussion below, we focus on the scaling relations in Eqs. (20) and (21).
@@ 4 Application of Scaling Relations to Individual Haloes and Comparison with other Approximations
We first show the $dM/dE$ and $f(E)$ constructed from simulations (referred to as “data”) for 8 representative haloes in Figs. 2 and 3, respectively. The $R_{\mathrm{vir}}$ and $M_{\mathrm{vir}}$ in the units for the axes of these figures are taken from the simulations for each halo. For comparison, we also show the $dM/dE$ and $f(E)$ calculated from the scaling relations in Eqs. (20) and (21), respectively, using the $r_{\mathrm{max}}(E)$ obtained from the simulated potential for each halo. For convenience, we refer to the latter results as obtained from “numerical scaling relations” (simplified to “scaling relations” when the context makes the meaning clear) because numerical potentials are used. Note that the $f(E)$ from the numerical scaling relation also uses the parameters $r_{\mathrm{s}}$ and $v_{\mathrm{s}}=r_{\mathrm{s}}\sqrt{4\pi G\rho_{\mathrm{s}}}$ of the best-fit NFW profile for each halo. To assess the optimal energy ranges for the scaling relations, we mark the numerical potentials at various radii by vertical lines in Figs. 2 and 3. It can be seen that the scaling relations are very good approximations to the data on individual haloes over broad energy ranges. For our entire sample of 79 haloes, we find that the average rms deviations are $\delta_{\log(dM/dE)}=0.04$ over $E\in[\Phi(0.02R_{\mathrm{vir}}),\Phi(R_{\mathrm{vir}})]$ and $\delta_{\log f(E)}=0.04$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(2.5R_{\mathrm{vir}})]$. These energy ranges correspond to the fitting ranges of the scaling relations. Because the common applicable range of the scaling relations are $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(R_{\mathrm{vir}})]$, we focus on $\delta_{\log(dM/dE)}$ and $\delta_{\log f(E)}$ over this range in comparing various approximations to the $dM/dE$ and $f(E)$ below (see Table 1). When appropriate, we also consider these deviations over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(2.5R_{\mathrm{vir}})]$.
@@ 4.2 Comparison of Results for NFW Profile
As discussed above, applying the numerical scaling relations requires $r_{\mathrm{max}}(E)$ from the simulated potential and the parameters of the best-fit NFW profile for a halo. It is interesting to examine the results from the scaling relations with $r_{\mathrm{max}}(E)$ from the potential of the best-fit NFW profile instead. We refer to these results as obtained from “NFW scaling relations” and show them for the 8 representative haloes in Figs. 2 and 3. For our entire sample of 79 haloes, we find that the average rms deviations are $\delta_{\log(dM/dE)}=0.06$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(R_{\mathrm{vir}})]$ and $\delta_{\log f(E)}=0.08$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(2.5R_{\mathrm{vir}})]$ (see Tabel 1). However, significant deviations occur for $E<\Phi(0.05R_{\mathrm{vir}})$ when the NFW potential differs significantly from the simulated one at $r<0.05R_{\mathrm{vir}}$ (see haloes with large values of $\Phi_{\rm diff}(0)=\Phi_{\rm NFW}(0)-\Phi(0)$ in Figs. 2 and 3).
As shown in Gross et al. (2024), the energy distribution and DF for a simulated halo can be described by those for the best-fit NFW profile to good approximation, but those NFW exact results have the same deficiency as the NFW scaling relations when the NFW potential differs significantly from the simulated one at $r<0.05R_{\mathrm{vir}}$. For our entire sample of 79 haloes, we find that the average rms deviations for the NFW exact results are $\delta_{\log(dM/dE)}=0.06$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(R_{\mathrm{vir}})]$ and $\delta_{\log f(E)}=0.12$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(2.5R_{\mathrm{vir}})]$ (see Tabel 1). So the NFW exact result for the DF is somewhat worse than the NFW scaling relation, while both approximations are comparable in the fit quality of the energy distribution (see Tabel 1).
@@ 5.1 Scaling Law for Volumetric Density of DF
Arad et al. (2004) found ${\cal{V}}(f)\propto f^{-2.50\pm 0.05}$ over a wide range of $f$ values for simulated halos. We now show that this power law can be accounted for by the scaling relations in Eqs. (20) and (21). For $f({\bm{r}},{\bm{v}})=f(E)$ and $df/dE<0$ (see Fig. 3), we rewrite Eq. (26) by a change of variable as
@@ 5.2 Origin of Scaling Relation for Energy Distribution
For an initial perturbation with $(\Delta M_{i}/M_{i})\propto M_{i}^{-\varepsilon}$, where $\varepsilon$ is determined by the index of the matter power spectrum on the relevant scale, a system under self-similar collapse in an Einstein-de Sitter Universe grows as $M_{{\mathrm{ta}}}(t)\propto t^{2/3\varepsilon}$ and $r_{{\mathrm{ta}}}(t)\propto M_{{\mathrm{ta}}}^{1/3}t^{2/3}\propto t^{(2+6\varepsilon)/9\varepsilon}$ (Mo et al., 2010). So we have $d\ln M_{{\mathrm{ta}}}(t_{\ast})/d\ln r_{{\mathrm{ta}}}(t_{\ast})=3/(1+3\varepsilon)$. Mass shells initially outside the target shell may move inside it during collapse. Due to this shell crossing and the resulting contraction of particle orbits, $M_{{\mathrm{ta}}}(t_{\ast})<M(r_{\mathrm{max}},t_{0})$ and $r_{{\mathrm{ta}}}(t_{\ast})>r_{\mathrm{max}}$. If the halo growth rate is low ($\varepsilon>2/3$) or if the accreted matter has non-negligible angular momentum, the inner density profile of a halo will soon settle down (Fillmore & Goldreich, 1984) so that $M(r_{\mathrm{max}},t)/M_{{\mathrm{ta}}}(t_{\ast})$ and $r_{\mathrm{max}}/r_{{\mathrm{ta}}}(t_{\ast})$ approach constants of order unity. Based on the above discussion, $\kappa$ is approximately a constant. We note that the above argument does not hold for fast, purely radial accretion ($\epsilon<2/3$). However, self-similar halo growth in this case leads to the SIS (Mo et al., 2010), for which a scaling relation [Eq. (14)] very close to that in Eq. (20) is obtained.
Neither self-similar halo growth with purely radial accretion nor the SIS with a divergent central potential are realistic models of haloes. Nevertheless, they represent reasonable approximations that can shed light on halo structure and formation. The existence of the scaling relation for $dM/dE$ in these simple models suggests that the scaling relations [Eqs. (20) and (21)] presented in this paper may be connected to the universal and fundamental properties imparted to haloes during their formation and relaxation.

### v2

@@ Abstract
pubyear: 2024pagerange: Scaling Relations in the Phase-Space Structure of Dark Matter Haloes5
@@ 1 Introduction
Other scaling relations for simulated haloes have been reported in the literature. Notably, as first presented by Taylor & Navarro (2001) and subsequently explored by many others (e.g., Austin et al. 2005; Hoffman et al. 2007; Ludlow et al. 2011; Arora & Williams 2020), the pseudo-phase-space density follows a power law, $\rho(r)/\sigma^{3}(r)\propto r^{-\gamma}$ with $\gamma=1.875$, where $\rho(r)$ and $\sigma(r)$ are the density and velocity dispersion at radius $r$, respectively. In addition, the volumetric density of the DF, ${\cal{V}}(f_{0})=\int d^{3}\bm{r}d^{3}\bm{v}\delta(f(\bm{r},\bm{v})-f_{0})$, was shown to approximately follow the power law ${\cal{V}}(f)\propto f^{-2.5}$ over a wide range of $f$ values (Arad et al., 2004). We will show that this result on ${\cal{V}}(f)$ can be accounted for by our scaling relations.
@@ 3 Scaling Relations for Simulated Haloes
Motivated by the SIS scaling relations, we now present similar scaling relations for simulated haloes. It is well known that the spherically-averaged density profile of dark matter haloes can be well fitted by the NFW profile (Navarro et al., 1997) between $0.05R_{\mathrm{vir}}$ and $R_{\mathrm{vir}}$ with $R_{\mathrm{vir}}$ being the virial radius, though deviations can occur outside this range (Schaller et al., 2015). The NFW profile is given by
@@
where $c=R_{\mathrm{vir}}/r_{\mathrm{s}}$ is the halo concentration, $\alpha$ and $\beta$ are dimensionless constants, and $m$ and $n$ are power indices. The NFW profile changes from $\rho\propto r^{-1}$ at $r\ll r_{\mathrm{s}}$ to $\rho\propto r^{-3}$ at $r\gg r_{\mathrm{s}}$, and behaves like the SIS at $r\sim r_{\mathrm{s}}$. While the SIS scaling relations suggest $m\approx 1$ and $n\approx-2$, we will obtain the best-fit values for $m$ and $n$ along with those for $\alpha$ and $\beta$ from the data on simulated haloes. The fitting procedure will also determine the ranges of $r_{\mathrm{max}}(E)$ for which the scaling relations provide good description of the data.
@@ 3.1 Halo Sample
As in Gross et al. (2024), we use the haloes in the TNG300-1-Dark simulations from the IllustrisTNG Project (Springel et al., 2017; Pillepich et al., 2017; Nelson et al., 2017; Naiman et al., 2018; Marinacci et al., 2018). The simulations were performed with the Planck Collaboration XIII (2016) cosmological parameters: $\Omega_{\mathrm{m}}=0.3089$, $\Omega_{\mathrm{b}}=0.0486$, $\Omega_{\Lambda}=0.6911$, $h=0.6774$, $n_{\mathrm{s}}=0.9667$, and $\sigma_{8}=0.8159$. The particle mass is $m_{\rm d}=7\times 10^{7}h^{-1}M_{\odot}$ and the softening length is $1h^{-1}\mathrm{kpc}$. The virial mass and radius of a halo, $M_{\mathrm{vir}}$ and $R_{\mathrm{vir}}$, respectively, are determined such that the average density inside $R_{\mathrm{vir}}$ is equal to $\Delta_{\mathrm{vir}}$ times the critical density of the Universe, where $\Delta_{\mathrm{vir}}=18\pi^{2}+82(\Omega_{\mathrm{m}}-1)-39(\Omega_{\mathrm{m}}-1)^{2}\approx 102$ is the virial factor (Bryan & Norman, 1998).
@@
For each halo of the final sample, we construct $dM/dE$ by counting the particles within $r_{\textrm{lim}}=2.5R_{\mathrm{vir}}$, which is approximately the depletion radius separating a growing halo from its draining environment (Fong & Han, 2021; Gao et al., 2023). As the energy is very sensitive to the choice of reference frame, we use the mean values for the most bound 10% of the particles in a halo to define its position and velocity (cf. Han et al., 2012). The DF is obtained as $f(E)=(dM/dE)/g(E)$ (e.g., Natarajan et al. 1997). Because we only count particles with $r<r_{\textrm{lim}}$, the corresponding $g(E)$ is
@@ 3.2 Fitting of Scaling Relations
Rather than fitting the scaling relation for $dM/dE$ in Eq. (17) to each halo, we aim to find a relation that applies to all the haloes. We obtain the numerical constant $\alpha$ and power index $m$ in this relation as follows. Because particles with $r\sim R_{\mathrm{vir}}$ are of main concern, we use a uniform logarithmic grid of $r_{\mathrm{max}}(E)/R_{\mathrm{vir}}$ for all the haloes. For each halo, we recalculate $dM/dE$ by counting the particles in each logarithmic bin of $r_{\mathrm{max}}(E)/R_{\mathrm{vir}}$ (i.e., $E$ is an implicit variable relating $r_{\mathrm{max}}(E)$ to $dM/dE$). Note that over the fitting range below, there are at least 10 particles in each bin for every halo so that $dM/dE$ can be calculated to reasonable accuracy. For each bin, we sort values of $(dM/dE)Gc^{1-m}/R_{\mathrm{vir}}$ for all the haloes to find the median. By inspection, the median value of $(dM/dE)Gc^{1-m}/R_{\mathrm{vir}}$ as a function of $r_{\mathrm{max}}(E)/R_{\mathrm{vir}}$ follows the scaling relation in Eq. (17) over $r_{\mathrm{max}}(E)\in[0.02R_{\mathrm{vir}},R_{\mathrm{vir}}]$ (this range is chosen to achieve the optimal fit). Detailed fitting by minimizing the rms deviation of $\log[(dM/dE)Gc^{1-m}/R_{\mathrm{vir}}]$ over this range gives $\alpha=0.60$ and $m=1.01$ with $\delta_{\log[(dM/dE)Gr_{\mathrm{s}}^{0.01}/R_{\mathrm{vir}}^{1.01}]}=0.019$. Because the fitted $m$ value is so close to 1, we repeat the above procedure by setting $m=1$ and obtain $\alpha=0.60$ with $\delta_{\log[(dM/dE)G/R_{\mathrm{vir}}]}=0.020$. In view of the simplicity without loss of accuracy, we adopt
@@
With $\delta_{\log[(dM/dE)G/R_{\mathrm{vir}}]}=0.019$ and $\delta_{\log[f(E)Gv_{\mathrm{s}}R_{\mathrm{vir}}^{2}]}=0.011$, these alternate scaling relations also provide good description of the halo data, but will not be discussed further in view of their similarity to the scaling relations adopted above.
We have also fitted the scaling relations in Eqs. (17) and (18) to individual haloes (see Appendix A). When uncertainties are taken into account, the sets of $(\alpha,m)$ and $(\beta,n)$ for individual haloes are consistent with those values fitted to the median halo results as shown in Eqs. (20) and (21). The intrinsic halo-to-halo scatter in these parameters is at the 1–3% level. For our discussion below, we focus on the scaling relations in Eqs. (20) and (21).
@@ 4 Application of Scaling Relations to Individual Haloes and Comparison with other Approximations
We first show the $dM/dE$ and $f(E)$ constructed from simulations (referred to as “data”) for 8 representative haloes in Figs. 2 and 3, respectively. The $R_{\mathrm{vir}}$ and $M_{\mathrm{vir}}$ in the units for the axes of these figures are taken from the simulations for each halo. For comparison, we also show the $dM/dE$ and $f(E)$ calculated from the scaling relations in Eqs. (20) and (21), respectively, using the $r_{\mathrm{max}}(E)$ obtained from the simulated potential for each halo. For convenience, we refer to the latter results as obtained from “numerical scaling relations” (simplified to “scaling relations” when the context makes the meaning clear) because numerical potentials are used. Note that the $f(E)$ from the numerical scaling relation also uses the parameters $r_{\mathrm{s}}$ and $v_{\mathrm{s}}=r_{\mathrm{s}}\sqrt{4\pi G\rho_{\mathrm{s}}}$ of the best-fit NFW profile for each halo. To assess the optimal energy ranges for the scaling relations, we mark the numerical potentials at various radii by vertical lines in Figs. 2 and 3. It can be seen that the scaling relations are very good approximations to the data on individual haloes over broad energy ranges [see also Appendix A, where the sets of $(\alpha,m)$ and $(\beta,n)$ for the scaling relations fitted to individual haloes are compared with those values used in Eqs. (20) and (21)]. For our entire sample of 79 haloes, we find that the average rms deviations are $\delta_{\log(dM/dE)}=0.04$ over $E\in[\Phi(0.02R_{\mathrm{vir}}),\Phi(R_{\mathrm{vir}})]$ and $\delta_{\log f(E)}=0.04$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(2.5R_{\mathrm{vir}})]$. These energy ranges correspond to the fitting ranges of the scaling relations. Because the common applicable range of the scaling relations are $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(R_{\mathrm{vir}})]$, we focus on $\delta_{\log(dM/dE)}$ and $\delta_{\log f(E)}$ over this range in comparing various approximations to the $dM/dE$ and $f(E)$ below (see Table 1). When appropriate, we also consider these deviations over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(2.5R_{\mathrm{vir}})]$.
@@ 4.2 Comparison of Results for NFW Profile
As discussed above, applying the numerical scaling relations requires $r_{\mathrm{max}}(E)$ from the simulated potential and the parameters of the best-fit NFW profile for a halo. It is interesting to examine the results from the scaling relations with $r_{\mathrm{max}}(E)$ from the potential of the best-fit NFW profile instead. We refer to these results as obtained from “NFW scaling relations” and show them for the 8 representative haloes in Figs. 2 and 3. For our entire sample of 79 haloes, we find that the average rms deviations are $\delta_{\log(dM/dE)}=0.06$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(R_{\mathrm{vir}})]$ and $\delta_{\log f(E)}=0.08$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(2.5R_{\mathrm{vir}})]$ (see Table 1). However, significant deviations occur for $E<\Phi(0.05R_{\mathrm{vir}})$ when the NFW potential differs significantly from the simulated one at $r<0.05R_{\mathrm{vir}}$ [see haloes with large values of $\Phi_{\rm diff}(0)=\Phi_{\rm NFW}(0)-\Phi(0)$ in Figs. 2 and 3]. Because $r_{\mathrm{max}}(E)$ is calculated from the potential, the difference in the potential between the NFW profile and the simulated halo results in different quality for the fits based on $r_{\mathrm{max}}(E)$.
As shown in Gross et al. (2024), the energy distribution and DF for a simulated halo can be described by those for the best-fit NFW profile to good approximation, but those NFW exact results have the same deficiency as the NFW scaling relations when the NFW potential differs significantly from the simulated one at $r<0.05R_{\mathrm{vir}}$. For our entire sample of 79 haloes, we find that the average rms deviations for the NFW exact results are $\delta_{\log(dM/dE)}=0.06$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(R_{\mathrm{vir}})]$ and $\delta_{\log f(E)}=0.12$ over $E\in[\Phi(0.05R_{\mathrm{vir}}),\Phi(2.5R_{\mathrm{vir}})]$ (see Table 1). So the NFW exact result for the DF is somewhat worse than the NFW scaling relation (see below), while both approximations are comparable in the fit quality of the energy distribution (see Table 1).
@@
As noted above, the NFW scaling relation gives a somewhat better fit to the DF than the NFW exact result (see Table  1). This improvement is mostly due to a better normalization coefficient incorporated in the fit to the simulated data. As shown in the right panel of Fig. 4, the NFW exact result for the DF lies consistently above the NFW scaling relation although there are only small differences in their shape over the $r_{\mathrm{max}}(E)$ range of interest.
@@ 5.1 Scaling Law for Volumetric Density of DF
Arad et al. (2004) found ${\cal{V}}(f)\propto f^{-2.50\pm 0.05}$ over a wide range of $f$ values for simulated haloes. We now show that this power law can be accounted for by the scaling relations in Eqs. (20) and (21). For $f({\bm{r}},{\bm{v}})=f(E)$ and $df/dE<0$ (see Fig. 3), we rewrite Eq. (26) by a change of variable as
@@ 5.2 Origin of Scaling Relation for Energy Distribution
For an initial perturbation with $(\Delta M_{i}/M_{i})\propto M_{i}^{-\varepsilon}$, where $\varepsilon$ is determined by the index of the matter power spectrum on the relevant scale, a system under self-similar collapse in an Einstein-de Sitter Universe grows as $M_{{\mathrm{ta}}}(t)\propto t^{2/3\varepsilon}$ and $r_{{\mathrm{ta}}}(t)\propto M_{{\mathrm{ta}}}^{1/3}t^{2/3}\propto t^{(2+6\varepsilon)/9\varepsilon}$ (Mo et al., 2010). So we have $d\ln M_{{\mathrm{ta}}}(t_{\ast})/d\ln r_{{\mathrm{ta}}}(t_{\ast})=3/(1+3\varepsilon)$. Mass shells initially outside the target shell may move inside it during collapse. Due to this shell crossing and the resulting contraction of particle orbits, $M_{{\mathrm{ta}}}(t_{\ast})<M(r_{\mathrm{max}},t_{0})$ and $r_{{\mathrm{ta}}}(t_{\ast})>r_{\mathrm{max}}$. If the halo growth rate is low ($\varepsilon>2/3$) or if the accreted matter has non-negligible angular momentum, the inner density profile of a halo will soon settle down (Fillmore & Goldreich, 1984, see also Zhao et al. 2003 for results of cosmological simulations) so that $M(r_{\mathrm{max}},t)/M_{{\mathrm{ta}}}(t_{\ast})$ and $r_{\mathrm{max}}/r_{{\mathrm{ta}}}(t_{\ast})$ approach constants of the order of unity, and ${d\ln r_{\mathrm{max}}}/{d\ln r_{{\mathrm{ta}}}(t_{\ast})}\simeq 1$. Based on the above discussion, $\kappa$ is approximately a constant. We note that the above argument does not hold for fast, purely radial accretion ($\epsilon<2/3$). However, self-similar halo growth in this case leads to the SIS (Mo et al., 2010), for which a scaling relation [Eq. (14)] very close to that in Eq. (20) is obtained.
Neither self-similar halo growth with purely radial accretion nor the SIS with a divergent central potential are realistic models of haloes. Nevertheless, they represent reasonable approximations that can shed light on halo structure and formation. The existence of the scaling relation for $dM/dE$ in these simple models suggests that the scaling relations [Eqs. (20) and (21)] presented in this paper may be connected to the universal and fundamental properties imparted to haloes during their formation and relaxation. A potential test of this scenario is the dependence of the coefficient $\kappa$ on the halo growth rate. For a higher accretion rate (i.e., a lower $\varepsilon$), we have a larger $d\ln M_{{\mathrm{ta}}}(t_{\ast})/d\ln r_{{\mathrm{ta}}}(t_{\ast})=3/(1+3\varepsilon)$. However, this trend is partly cancelled by a smaller $M_{{\mathrm{ta}}}(t_{\ast})/M(r_{\mathrm{max}},t_{0})$ due to stronger shell crossing of recent accretion (Fillmore & Goldreich 1984; Shi 2016). Therefore, the dependence of $\kappa$ on the halo growth rate is weak, which is consistent with the relatively small scatter in the set of $(\alpha,m)$ for the scaling relation fitted to the $dM/dE$ for individual haloes of different growth histories (see Appendix A). This approximate universality makes the scaling relations even more interesting, and also adds to the challenge of understanding their origin.